OpenStudy (anonymous):

Using the completing-the-square method, rewrite f(x) = x2 − 8x + 3 in vertex form.

OpenStudy (zzr0ck3r):

\(\large x^2+bx+c=(x+\frac{b}{2})^2+c-\frac{b^2}{4}\)

OpenStudy (jhannybean):

I'll show you how @zzr0ck3r derived this awesome formula! \[x^2-8x+3\]\[(x^2-8x)+3\] \[c=\left(-\frac{8}{2}\right)^2 = (-4)^2 = 16\] \[(x^2-8x+16)+3-16\]\[\boxed{(x-4)^2-13} \iff (x+\frac{b}{2})^2+c-\frac{b^2}{4}\]

OpenStudy (jhannybean):

Maybe not derived... not the right word... but how your function correlates to the formula*